OpenStudy (anonymous):

What is one twelifth of 1889332?

OpenStudy (turingtest):

what is one-twelfth written as a fraction with numbers?

OpenStudy (anonymous):

Yes, What would it be if you took one twelth out of that number?

OpenStudy (turingtest):

you did not answer my question :(

OpenStudy (anonymous):

Just written as numbers.

OpenStudy (turingtest):

how do you write "one-twelfth" written as a fraction (with numbers)? e.g. one-half written as a fraction is 1/2 so what is one-twelfth?

OpenStudy (anonymous):

You don't write it as a fraction with numbers you would round it once you take it out with your left out numbers of your fraction

OpenStudy (cwrw238):

use your calculator or wolfram

OpenStudy (turingtest):

are you telling me that I am not taking the right approach to this problem? I assure you if you answer my question you will get your answer.

OpenStudy (anonymous):

Wolfram said it would be 7 decimal digits.

OpenStudy (turingtest):

?

OpenStudy (anonymous):

This is excatley what the question says "One Twelfth of 1189332?" and it's on a Crossword Puzzel.

OpenStudy (turingtest):

post me the link you got from wolfram so I can see what you are doing. I don't know about the puzzle, I'm just answering the math part...

OpenStudy (anonymous):

This is the excat link: http://www.wolframalpha.com/input/?i=What+is+one+twelifth+of+1889332%3F

OpenStudy (turingtest):

ok, you don;t know how to use wolfram... so let's go back to my question: how do you write "one-twelfth" written as a fraction (with numbers)? e.g. one-half written as a fraction is 1/2 so what is one-twelfth?

OpenStudy (anonymous):

1.12

OpenStudy (anonymous):

I think

OpenStudy (turingtest):

no one-half=1/2 one-third=1/3 one-tenth=1/10 one-twelfth=?

OpenStudy (anonymous):

One-Tweflth would be 0.83 as a decimal and as a fraction 1/144

OpenStudy (turingtest):

how did you get that o-0 the decimal is almost right...

OpenStudy (turingtest):

did you put something in your calculator? if so, what?

OpenStudy (turingtest):

forget it... here's the pattern: one-half=1/2 one-fifth=1/5 one-tenth=1/10 one-eleventh=1/11 so... one-twelfth=1/12 does this make sense now?

OpenStudy (anonymous):

Yes now it does

OpenStudy (turingtest):

ok, now I will tell you that the word "of" in mathematics translates to "times" so what is one-twelfth of 1889332

OpenStudy (anonymous):

he put into calculator 1/12 and decimal 0.8333333333 and then squared it, lol

OpenStudy (turingtest):

goodness... Trying to be patient...

OpenStudy (anonymous):

Yeah that was what I did but thats the reason why I don't know. I know one-twelfth as a fraction would be 1/12 so then you would be dividing 1889332?

OpenStudy (turingtest):

Yes! dividing it by what?

OpenStudy (anonymous):

1889332?

OpenStudy (anonymous):

Lol I'm sorry but Math is not my best subject.

OpenStudy (turingtest):

one-twelfth=1/12 of=times what is What is one-twelfth of 1889332? becomes... what is \[\frac{1}{12}\times1889332\]

OpenStudy (anonymous):

I have no clue I have tried.

OpenStudy (turingtest):

use a calculator...

OpenStudy (anonymous):

Just tell all of my calcutors don't have abc key.

OpenStudy (turingtest):

what if I tell you that \[\frac{1}{12}\times1889332=\frac{1889332}{12}\]can your calculator do that?

OpenStudy (anonymous):

472033/2

OpenStudy (anonymous):

W

OpenStudy (anonymous):

We have been here an Hour just for one problem that I need but apprently you don't want to move on to your next person you have to help lol.

OpenStudy (turingtest):

Wow, I have no idea where you got 472033/2 look, I'm not a free-answer person, but here I will make an exception because I am tired \[\frac{1889332}{12}=\frac{472333}{3}=57444.333...\]that is all g'night, and please study.

OpenStudy (anonymous):

for one twelfth , you divide by 12

OpenStudy (turingtest):

yes you do...

OpenStudy (anonymous):

haha finally u finished :D